OpenStudy (anonymous):

Write the following equation in q(x)+(r(x)/d(x)) form: (7x^3 - 11x^2 + 5x +4) / (x-1)

OpenStudy (helder_edwin):

using synthetic divison u get: \[\large \begin{array}{r|rrrr} & 7 & -11 & 5 & 4\\ 1 & & 7 & -4 & 1\\ \hline & 7 & -4 & 1 & \fbox{5} \end{array} \]

OpenStudy (anonymous):

okay but once I use synthetic division how do I put it in the form its asking for? @helder_edwin

OpenStudy (anonymous):

and where did you get your second row of numbers for the synthetic division?

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

1*7=7 -11+7=-4 1*-4=-4 5+(-4)=5-4=1 1*1=1 4+1=5 \[q(x)=7 x^2-4 x+1,r(x)=5\] d(x)=x-1

OpenStudy (anonymous):

oh! ok awesome thank you so much!!! I get it now!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

okay so its (7x^2 - 4x + 1)+(5 / (x - 1)) ??? right?

OpenStudy (anonymous):

correct. \[\frac{ 7x^3-11x^2+5x+4 }{ x-1 }=7x^2-4x+1+\frac{ 5 }{ x-1 }\]

OpenStudy (anonymous):

and hey for example if it was divided by (x-2) instead would it look like this?: 

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@mathlover2014

OpenStudy (anonymous):